I have 20 more nickels than i have dimes if the sum of the money i have is $8.50 how many coins of each type do I have?
Answer by macston(5194)


N=no. of nickels; D=no. of dimes=N-20
$0.05N + $0.10D =$8.50 Substitute for D
$0.05N+$0.10(N-20)=$8.50
$0.05N+$0.10N-$2.00=$8.50 Add $2.00 to each side.
$0.15N=$10.50 Divide each side by $0.15.
N=70 ANSWER 1:You have 70 nickels
D=N-20=70-20=50 ANSWER 2; You have 50 dimes.
CHECK
$0.05N + $0.10D =$8.50
$0.05(70) + $0.10(50) =$8.50
$3.50+$5.00=$8.50
$8.50=$8.50